# Polarization modes of gravitational waves in general symmetric teleparallel gravity

### v1

@@ IV.3 Scalar modes
There are six scalar equations, which are relatively complex, so we have placed the detailed forms of the equations in Appendix A. The scalar polarization modes depend on the choice of parameters, and their general analysis is very complex and tedious. However, two universal conclusions can still be extracted. The first conclusion is that there always exists the following plane wave solution propagating at the speed of light: $L+K=0,\phi=\Theta=0$. Therefore, according to Eq. (III), in the case where the test particles have hypermomentum, there always exists a longitudinal mode propagating at the speed of light. The second conclusion is that, given the homogeneity of the field equations and their formal adherence to Lorentz symmetry, any possible scalar modes must propagate at the speed of light.
@@ V.1 $f(Q)$ gravity
$f(Q)=\alpha Q+\beta Q^{2}+...$ (43)
It can be seen that the parameter corresponding to $f(Q)$ theory is
@@
Therefore, in the case where test particles do not have hypermomentum, $f(Q)$ theory only has tensor modes ($+$ and $\times$ modes) propagating at the speed of light, without additional vector or scalar modes, which is consistent with the conclusion in Ref. S Capozziello . In the case with hypermomentum charge, according to Eq. (III), we find that, in addition to the tensor modes propagating at the speed of light, the theory always has unconstrained shear modes and the unconstrained longitudinal mode. This is physically unreasonable, and therefore, a viable $f(Q)$ theory must a priori assume that the matter field action is independent of the connection.
However, under the assumption of no hypermomentum, Refs. D. Aguiar Gomes ; Lavinia Heisenberg6 studied cosmological perturbations in $f(Q)$ theory with the typical scalar matter field $S_{m}=\int d^{4}x\left[-(\partial\Phi)^{2}-V(\Phi)\right]$ and found that ghost and strong coupling problems are widespread in $f(Q)$ gravity. Reference Lavinia Heisenberg6 suggests that coupling the connection with matter fields could be a potential solution to avoid these problems. However, our study demonstrates that this approach leads to observable pathological polarization modes of gravitational waves, effectively ruling out this possibility. In conclusion, whether or not hypermomentum is assumed, pathological issues typically arise in $f(Q)$ theory. Our study further highlights the importance of exercising caution when working with this theory.
@@ V.2 Quadratic non-metricity gravity
The parameters corresponding to this theory are as follows:
@@
and all other parameters are zero. In this case, scalar polarization modes of gravitational waves still depend on the parameters.
In the literature on quadratic non-metricity gravity, the following two parameter conditions are frequently mentioned Lavinia Heisenberg2 ; Lavinia Heisenberg3 ; Jose Beltran Jimenez :
Parameter (a)
@@
By imposing the coincident gauge $\Sigma^{\lambda}_{~{}\mu\nu}=0$ at the level of the second-order perturbation action while retaining only the variable $h_{\mu\nu}$, these two parameter conditions respectively endow the second-order perturbation action with the gauge symmetry and the WTDiff symmetry Lavinia Heisenberg2 ; Lavinia Heisenberg3 ; Jose Beltran Jimenez in form. We now analyze the polarization modes of gravitational waves under these two parameters.
For Parameter (a), we note that in this case, $C_{(1)}+E_{(1)}=4c_{1}+2c_{2}+2c_{4}=0$. Therefore, there must be an unconstrained $K_{i}$ in the theory, leading to a pathological situation with hypermomentum. For the case without hypermomentum, there are no vector mode gravitational waves. For scalar modes, it gives the solution $\phi=\Theta=0$, while imposing no constraints on $K$ and $L$. Therefore, in the absence of hypermomentum, scalar mode gravitational waves do not exist. In conclusion, the theory only has $+$ and $\times$ modes that propagate at the speed of light. For Parameter (b), the same reasoning leads to the conclusion that only the case with no hypermomentum can be considered reasonable. Similarly, in the absence of hypermomentum charge, there are no vector mode gravitational waves. In this case, there is no scalar mode, although $L$ and $K$ can propagate. In this parameter case, only $+$ and $\times$ modes propagating at the speed of light exist.
Regarding quadratic non-metricity gravity, one additional point needs to be emphasized. It can be seen that there are six free parameters $A_{(1)},B_{(1)},C_{(1)},D_{(1)},E_{(1)}$, and $A_{(2)}$ contributing to the linear field equations of general symmetry teleparallel gravity, while quadratic non-metricity gravity only has five free parameters. This indicates that even at the linear order, there are theories that can not be encompassed by quadratic non-metricity gravity, which includes all possible quadratic terms of $Q_{\lambda\mu\nu}$. However, we note that the following action can encompass all six free parameters:
@@
Some may think that symmetry teleparallel gravity requires the curvature to be zero, so $R$ does not contribute to the action. However, it is important to emphasize that the condition of zero curvature is imposed at the level of the field equations, not at the level of the action. Therefore, one cannot naively impose constraints in the action, but rather must consider and prove this more carefully, or else errors may arise.
@@ VI Conclusion
In this paper, we studied the polarization modes of gravitational waves in the most general symmetric teleparallel gravity, where second-order field equations can be derived, both with and without hypermomentum. We found that, in the case without hypermomentum, the theory always contains two tensor polarization modes of gravitational waves, the $+$ and $\times$ modes, propagating at the speed of light. Unless in an extremely special parameter case: $04A_{(2)}+C_{(1)}+E_{(1)}=$ 0, the theory generally does not have vector modes. Scalar modes depend on the choice of parameters, but the possible scalar gravitational waves can only propagate at the speed of light. In the case with hypermomentum, there are always $+$ and $\times$ modes propagating at the speed of light. For vector modes, either there exist shear modes that propagate at the speed of light, or the shear modes are pathological. The vector-$x$ and vector-$y$ modes also exist only under a very specific parameter condition $04A_{(2)}+C_{(1)}+E_{(1)}=$ 0. Scalar modes also depend on the parameters, but there is always a longitudinal mode that propagates at the speed of light. Similarly, the possible scalar mode must propagate at the speed of light. We found that, in the presence of hypermomentum, the additional shear modes and the longitudinal mode are universally present. This characteristic can therefore serve as an important criterion for distinguishing symmetric teleparallel gravity from other theories.
@@
We also specifically considered two typical symmetric teleparallel gravity theories: $f(Q)$ gravity and quadratic non-metricity gravity. For $f(Q)$ gravity, we found that assuming any coupling between the matter field and the connection, i.e., the presence of hypermomentum, leads to this theory being physically unreasonable. Therefore, within the framework of $f(Q)$ theory, it can only be assumed that the matter field couples solely to the metric, and in this case, the theory has only two tensor modes that propagate at the speed of light. Combining the analysis of the ghost and strong coupling problems in the context of cosmological perturbations from Refs. D. Aguiar Gomes ; Lavinia Heisenberg6 , our results further suggest that $f(Q)$ theory must be treated with caution, as it may exhibit pathological behavior.
@@
For quadratic non-metricity gravity, we performed a detailed analysis of its gravitational wave polarization modes under two typical specific parameters. We found that under both sets of parameters, the assumption of the existence of hypermomentum can not be made. In the absence of hypermomentum, only tensor modes propagating at the speed of light exist under both parameter sets. We also pointed out that, at the linear level, quadratic non-metricity gravity can not cover the most general linearized field equations. By introducing the scalar curvature $R$ into the action, we found a class of theories that can cover the most linearized field equations at the linear level. The nontrivial contribution of $R$ may come as a surprise to some.
@@ Acknowledgments
This work is supported in part by the National Natural Science Foundation of China (Grants No. 123B2074, No. 12475056, and No. 12247101), Gansu Province’s Top Leading Talent Support Plan, the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06), the Natural Science Foundation of Gansu Province (No. 22JR5RA389), and the ‘111 Center’ under Grant No. B20063.
@@ Appendix A The specific forms of scalar equations
$\left(-\frac{1}{2}A_{(1)}-\frac{1}{4}C_{(1)}-\frac{3}{4}E_{(1)}-B_{(1)}\right)\partial_{0}K+\left(\frac{1}{4}C_{(1)}+\frac{1}{4}E_{(1)}\right)\frac{\Delta}{\partial_{0}}K$
@@
$-\left(\frac{3}{4}C_{(1)}+\frac{3}{4}E_{(1)}\right)\frac{\partial_{0}^{3}}{\Delta}\Theta+\left(4A_{(2)}+\frac{7}{4}C_{(1)}+\frac{7}{4}E_{(1)}\right)\partial_{0}\Theta=0,$ (55)

### v2

@@ IV.3 Scalar modes
There are six scalar equations, which are relatively complex, so we have placed the detailed forms of these equations in Appendix A. The properties of these scalar polarization modes depend on the parameter space, and their general analysis is very complex and tedious. However, two universal conclusions can still be extracted. The first one is that there always exists the following plane wave solution propagating at the speed of light: $L+K=0,\phi=\Theta=0$. Therefore, according to Eq. (III), in the case where the test particles have hypermomentum, there always exists a longitudinal mode propagating at the speed of light. The second one is that, owing to the homogeneity of the field equations and their formal Lorentz symmetry, all possible scalar modes must propagate at the speed of light.
@@ V.1 $f(Q)$ gravity
$f(Q)=\alpha Q+\beta Q^{2}+....$ (43)
It can be seen that the parameters corresponding to $f(Q)$ theory are
@@
Therefore, in the case where test particles do not have hypermomentum, $f(Q)$ theory only has tensor modes ($+$ and $\times$ modes) propagating at the speed of light, without additional vector or scalar modes, which is consistent with the conclusion obtained in Ref. S Capozziello . In the case with hypermomentum charge, according to Eq. (III), we find that, in addition to the tensor modes propagating at the speed of light, the theory always has unconstrained shear modes and the unconstrained longitudinal mode. This is physically unreasonable, and therefore, a viable $f(Q)$ theory must a priori assume that the matter field action is independent of the connection.
However, assuming the absence of hypermomentum, Refs. D. Aguiar Gomes ; Lavinia Heisenberg6 studied cosmological perturbations in $f(Q)$ theory with the typical scalar matter field $S_{m}=\int d^{4}x\left[-(\partial\Phi)^{2}-V(\Phi)\right]$ and found that ghost and strong coupling problems are widespread in $f(Q)$ gravity. It was suggested in Ref. Lavinia Heisenberg6 that introducing couplings between the connection and matter fields could be a potential solution to avoid these problems. However, our study demonstrates that this approach leads to observable pathological polarization modes of gravitational waves, effectively ruling out this possibility. In conclusion, whether or not hypermomentum is assumed, pathological issues typically arise in $f(Q)$ theory. Our study further highlights the importance of exercising caution when working with this theory.
@@ V.2 Quadratic non-metricity gravity
References Ismail Soudi ; Manuel Hohmann have already conducted a certain level of investigation into the gravitational wave polarization modes of this theory in the absence of hypermomentum. The parameters corresponding to this theory are as follows:
@@
and all other parameters are zero. In this case, scalar polarization modes of gravitational waves still depend on the parameter space.
In the literature on quadratic non-metricity gravity, the following two conditions for these parameters are frequently mentioned Lavinia Heisenberg2 ; Lavinia Heisenberg3 ; Jose Beltran Jimenez :
carameter (a)
@@
By imposing the coincident gauge $\Sigma^{\lambda}_{~{}\mu\nu}=0$ at the level of the second-order perturbation action while retaining only the variable $h_{\mu\nu}$, the above two conditions respectively endow the second-order perturbation action with the gauge symmetry and the Weyl Transverse Diffeomorphism (WTDiff) symmetry Lavinia Heisenberg2 ; Lavinia Heisenberg3 ; Jose Beltran Jimenez in form. We now analyze the polarization modes of gravitational waves under these two conditions.
For condition (a), we have $C_{(1)}+E_{(1)}=4c_{1}+2c_{2}+2c_{4}=0$. Therefore, there must be an unconstrained $K_{i}$ in the theory, leading to a pathological situation with hypermomentum. For the case without hypermomentum, there are no vector-mode gravitational waves. For scalar modes, it gives the solution $\phi=\Theta=0$, while imposing no constraints on $K$ and $L$. Therefore, in the absence of hypermomentum, scalar-mode gravitational waves do not exist. In conclusion, the theory only includes $+$ and $\times$ modes that propagate at the speed of light. For condition (b), the same reasoning leads to the conclusion that only the case with no hypermomentum is reasonable. Similarly, in the absence of hypermomentum charge, there are no vector-mode gravitational waves. In this case, there is no scalar mode, although $L$ and $K$ can propagate. In this condition, this theory predicts only $+$ and $\times$ modes propagating at the speed of light.
Regarding quadratic non-metricity gravity, one additional point needs to be emphasized. It is evident that the linear field equations of general symmetric teleparallel gravity involve six free parameters—$A_{(1)},B_{(1)},C_{(1)},D_{(1)},E_{(1)}$, and $A_{(2)}$—whereas quadratic non-metricity gravity contains only five independent parameters. This indicates that even at the linear order, there are theories that can not be encompassed by quadratic non-metricity gravity, which includes all possible quadratic terms of $Q_{\lambda\mu\nu}$. However, we note that the following action can encompass all six free parameters:
@@
A view holds that symmetry teleparallel gravity requires the curvature to be zero, so $R$ does not contribute to the action. However, it is important to emphasize that the condition of zero curvature is imposed at the level of the field equations, not at the level of the action. Therefore, one cannot naively impose constraints on the action, but rather must carefully consider and justify them; otherwise, errors may arise.
@@ VI Conclusion
In this paper, we studied the polarization modes of gravitational waves in the most general symmetric teleparallel gravity, where second-order field equations can be derived, both with and without hypermomentum. We found that, in the case without hypermomentum, the theory always contains two tensor polarization modes of gravitational waves, the $+$ and $\times$ modes, propagating at the speed of light. Unless in an extremely special condition for the parameters: $04A_{(2)}+C_{(1)}+E_{(1)}=$ 0, the theory generally does not have vector modes. The existence of scalar modes depends on the choice of parameters, but the possible scalar gravitational waves can only propagate at the speed of light. In the case with hypermomentum, there are always two tensor modes propagating at the speed of light. For vector modes, shear modes must either propagate at the speed of light or be pathological (they are not constrained by the linearized field equations). The vector-$x$ and vector-$y$ modes are only permissible under the specific parameter constraint $04A_{(2)}+C_{(1)}+E_{(1)}=$ 0. While scalar modes similarly depend on system parameters, but there always exists a longitudinal mode that propagates at the speed of light regardless of parameter values. Our analysis reveals that, in the presence of hypermomentum, the additional shear modes and the longitudinal mode are universally present. This distinctive characteristic provides an important criterion for distinguishing symmetric teleparallel gravity from other gravitational theories.
@@
We also specifically considered two typical symmetric teleparallel gravity theories: $f(Q)$ gravity and quadratic non-metricity gravity. For $f(Q)$ gravity, we found that assuming any coupling between the matter field and the connection, i.e., the presence of hypermomentum, renders the theory physically unreasonable. Therefore, within the framework of $f(Q)$ theory, the matter field must couple solely to the metric, and in this case, the theory has only two tensor modes that propagate at the speed of light. Combining the analysis of the ghost and strong coupling problems in the context of cosmological perturbations from Refs. D. Aguiar Gomes ; Lavinia Heisenberg6 , our results further suggest that $f(Q)$ theory must be treated with caution, as it may exhibit pathological behavior.
@@
For quadratic non-metricity gravity, we performed a detailed analysis of its gravitational wave polarization modes under two specific parameter conditions. We found that under both conditions, the assumption of the existence of hypermomentum can not be made. In the absence of hypermomentum, only tensor modes propagating at the speed of light exist. We also pointed out that, at the linear level, quadratic non-metricity gravity can not cover the most general linearized field equations. By introducing the scalar curvature $R$ into the action, we found a class of theories that can cover the most linearized field equations at the linear level. The nontrivial contribution of $R$ may lead to unexpected results.
@@ Acknowledgments
This work is supported in part by the National Key Research and Development Program of China (GrantNo. 2020YFC2201503), the National Natural Science Foundation of China (Grants No. 123B2074, No. 12475056, and No. 12247101), Gansu Province’s Top Leading Talent Support Plan, the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2024-jdzx06), the Natural Science Foundation of Gansu Province (No. 22JR5RA389), and the ‘111 Center’ under Grant No. B20063.
@@ Appendix A The specific forms of scalar equations
$\left(-\frac{1}{2}A_{(1)}-\frac{1}{4}C_{(1)}-\frac{3}{4}E_{(1)}-B_{(1)}\right)\partial_{0}K+\left(\frac{1}{4}C_{(1)}+\frac{1}{4}E_{(1)}\right)\frac{\Delta}{\partial_{0}}K$ (55)
@@
$-\left(\frac{3}{4}C_{(1)}+\frac{3}{4}E_{(1)}\right)\frac{\partial_{0}^{3}}{\Delta}\Theta+\left(4A_{(2)}+\frac{7}{4}C_{(1)}+\frac{7}{4}E_{(1)}\right)\partial_{0}\Theta=0,$